Installation of MIFAYN

STRASSER Michel December 18, 2001

Contents

I Quick installation 2

1 Unpacking 2 2 First upgrade 2 3 config.php3 3 4 The key 3 5 The end 4 6 Editing the config file with a spreadsheet 4

II Extending Mifayn 5

Part I

Quick installation

1 Unpacking

If you have GNU tar, simply do :

tar xzvf mifayn1.0.0.tgz

else do

gunzip mifayn1.0.0.tgz
tar xvf mifayn1.0.0

Move the src directory in a subtree of your exported web directory :

mv mifayn1.0.0/src/ /var/www/html/mysite/

for example ...

Changing some permissions (mifayn must create files):

chmod 777 /var/www/html/mysite/
chmod 777 /var/www/html/mysite/upgrade.htm
chmod 777 /var/www/html/mysite/img/

2 First upgrade

Before changing our configuration file, we are checking that everythings go right with the ”packaged” config file.

  Open a web browser, make it point to http://127.0.0.1/mysite/upgrade.

htm. In the Mot de passe / Passwort text zone, you should enter ”love” which is the current password. In the Fichier de configuration text zone, you should browse for the v 001.csv file which is in /mifayn1.0.0/config file .

Then press Envoyer / schicken .

If everything works well you should obtain figure 1 .

If something works wrong, Mifayn should tell you the nature of the error.

Figure 1: SuccedingCompilation

As the configuration file has been tested, an error at this step can occurs if you haven’t set the right permissions for the Mifayn directory. Mifayn wants to create files, you must allow him.

Some errors can happen due to your php installation that might be incompatible with this soft. (please contact me in this case).

If things works well, new files have been generated. Make a tour by following the links.

3 config.php3

Some variables must be set in the config.php3 file (in /var/www/html/mysite/ ). Edit this file. Change the password. Change the key. Edit the configuration file( v 001.csv ) (with a spreadsheet (see section 6) or with an editor), copy / paste the key that you have changed in the config file. Change the website name.

4 The key

The key is a protection mecanism that protects you from people who wants to change the content of your website.

5 The end

Your web site is now installed, you can customize it by editing the config file. First read the sample config file and try to understand.

6 Editing the config file with a spreadsheet

You can edit the configuration file with a spreadsheet. Import the file and set that the separator character is TAB and that lines are UNIX formatted. Read the User Manual for more informations.

Part II

Extending Mifayn

You can extend Mifayn by writing themes or builders, I have to write this section later.